# D4C: Improving Negative Example Quality to Enhance Machine Abstract Reasoning Ability

### v14

@@ IX-A Experiment on NICO
D3C enhances the positive-negative example contrast process in traditional visual deep learning models and thereby improves the quality of negative examples. Therefore, to attach D4C to these visual models, D3C must serve as the foundation. When using the D4C method to provide high-quality negative examples for the visual model processing NICO data, we maintain a memory bank for each category in the NICO dataset, that is, we maintain an $M_{n-1}$ and $Z_{n-1}$ for each category. When generating high-quality negative example representations for a specific subject category in NICO, we input the sample representations of other categories into the negative example generator. Then, the corresponding memory banks $M_{n-1}$ and $Z_{n-1}$ are selected based on the subject category to assist in the generation process. The experimental results are recorded in Table IV.
@@
TABLE V: various net on NICO 8:2
@@ IX-B Experiment on RAVEN
We tested Lico-Net on the RAVEN and I-RAVEN datasets to showcase its potential. In our exploration of the maximum potential of Lico-Net in solving RPM problems, we decided to discard the use of Sinkhorn distance for de-fuzzification and instead adopted a Transformer-Encoder equipped with learnable class vectors. Initially, we employed the Sinkhorn distance in preliminary experiments to investigate how Lico-Net induces RPM concepts. However, in order to accelerate model convergence and enhance its flexibility, we have chosen the learnable module of Transformer-Encoder to undertake the task of de-fuzzification.
Additionally, we conducted experiments combining D3C and D4C with Lico-Net to showcase the effectiveness of the D3C and D4C methods. For a fair comparison, we used the same settings and equipment as the current state-of-the-art (SOTA) method, Triple-CFN[37], including data volume, optimizer parameters, data augmentation, batch size, and other hyperparameters.
@@
An interesting observation from the Table VI indicates a commonality among existing excellent RPM inference engines, including Lico-Net. Specifically, these inference engines exhibit varying reasoning accuracy when tackling the RAVEN [17] and I-RAVEN [40] datasets. The I-RAVEN dataset was designed to address certain limitations in the option pool configuration of RAVEN by introducing improvements to its negative example setup. These enhancements include, but are not limited to, increasing the number of attributes that are altered in the negative examples and augmenting the magnitude of attribute deviations within them. However, from a human perspective, since the setup of positive examples remains unchanged, such additional setups do not alter the difficulty of reasoning, making I-RAVEN neither more nor less challenging than RAVEN. This paper speculates that the disparate outcomes achieved by these models on RAVEN and I-RAVEN could potentially be attributed to the higher quality of negative examples present in the I-RAVEN dataset compared to those in RAVEN. This phenomenon, to some extent, also supports the analysis presented in this paper.
The Table VI also conveys that Lico-Net demonstrates impressive performance as an Inference engine. Furthermore, D4C not only enhances Lico-Net’s performance but also, to some extent, bridges the gap in Lico-Net’s performance between the RAVEN and I-RAVEN datasets.
@@ IX-C Experiment on Bongard-Logo
We conducted experiments using Lico-Net-Bongard on the Bongard-Logo dataset; additionally, we explored the enhancement effects of the D3C and D4C algorithms on a series of excellent Bongard-Logo reasoning engines, including Lico-Net-Bongard. The experimental results are presented in the Table VII.
@@
It is noteworthy that, through experiments, we found that the negative example generator proposed in the D4C algorithm excels at generating fuzzy representations of high-scoring negative examples. This often leads to the generator easily causing gradient explosion during the training of Triple-CFN and our designed Lico-Net-Bongard. The solution presented in this paper is to perform a weighted summation of the generated negative example representations and the representations of the negative examples provided in the dataset, with a weight ratio of 0.2 to 0.8.
Finally, this paper notes that augmentation experiments were conducted for both PMoC [36] and Triple-CFN [37]. These experiments involved expanding the data volume of the Bongard-Logo dataset using data augmentation techniques, specifically including rotation and flipping. It should be emphasized that these augmentation techniques do not explicitly alter the diversity of concepts or the proportion between concepts inherent in the Bongard-Logo dataset [36]. Rather, the aim of these experiments was to investigate the potential of the models with a larger dataset. To verify the effectiveness of D4C on the augmented dataset, we conducted experiments, and the results are presented in Table VIII.
@@
Surprisingly, Table VIII demonstrates that Lico-Net-Bongard exhibits superior performance on the augmented Bongard-Logo problem, suggesting that Lico-Net-Bongard, designed based on fuzzy system methodology [14, 15], may possess greater network capacity and stronger learning abilities. Especially when combined with the D4C algorithm, Lico-Net-Bongard’s performance on the augmented Bongard-Logo problem is particularly outstanding, fully proving the effectiveness and potential of fuzzy systems as a methodological guide for deep learning algorithms.
@@ IX-D Experiment on PGM
In this paper, we conducted experiments related to Lico-Net on PGM. And the results are recorded in Tabel IX. The data recorded in the table reveals many meaningful phenomena.
@@
The combinations of Lico-Net+D3C, Triple-CFN+D3C, and RS-Tran+D3C exhibited very limited improvement in reasoning accuracy, which was expected. After all, there are fundamental differences between RPM and Bongard-Logo. RPM problems share a common concept space and attribute recognition mechanism, while D3C, designed to address concept interference and erroneous coupling issues between different samples, is less effective for RPM problems, as anticipated. In contrast, D4C brought significant performance improvements to Lico-Net, Triple-CFN, and RS-Tran. This fully validates our hypothesis that enhancing the quality of negative samples can effectively boost the performance of deep learning algorithms. This finding holds important theoretical and practical significance.
Additionally, we conducted experiments on the generalization issue in PGM for Lico-Net, D3C, and D4C. The experimental results are recorded in Table X.
@@ IX-E Ablation experiments
We removed the linear attention mechanism from the D4C generator while retaining the process of the transformer-encoder for processing input representations, and conducted relevant experiments again. The experimental results are recorded in the entry labeled “D4C-” in the Table XI, XII and XIII.
@@
It can be observed from the experimental results recored in these three tables that the step-wise linear attention mechanism serves as a crucial external memory bank for the D4C negative example generator, providing stable performance improvements.
@@ X Conclusion
This paper points out that when deep learning is applied to problems that embed complex and deep human concepts, deep learning algorithms should explicitly extract these concepts as fuzzy representations and reason based on these representations. This is also a key methodology in fuzzy systems. The Lico-Net and Lico-Net-Bongard designed in this paper both demonstrate the validity of this viewpoint and prove that fuzzy systems should be one of the key development directions for deep learning.
More notably, this paper proposes that deep learning models tasked with discriminative tasks are more concerned with the distinctions between positive and negative examples rather than the logic of discriminating between them. Focusing on the distinctions between positive and negative examples is feasible and effective in traditional discriminative tasks. However, when the quality of negative examples is not high enough to strictly define the boundaries of concepts, the performance of the algorithm may be compromised. Therefore, this paper suggests that providing stricter negative examples for deep learning models can effectively improve their performance. Consequently, this paper designs the GSCLM algorithm for NICO data and a negative example generator based on a step-wise linear attention mechanism for RPM problems, Bongard-Logo problems, and NICO problems. The results of the experiment demonstrate the effectiveness of these algorithms. Specifically, this generator represents a novel attempt within the methodology of fuzzy systems, and its effectiveness hints at the potential of fuzzy systems in the field of artificial intelligence.

### v16

@@ IX-A Experiment on NICO
D3C enhances the positive-negative example contrast process in traditional visual deep learning models and thereby improves the quality of negative examples. Therefore, to attach D4C to these visual models, D3C must serve as the foundation. When using the D4C method to provide high-quality negative examples for the visual model processing NICO data, we maintain a memory bank for each category in the NICO dataset, that is, we maintain an $M_{n-1}$ and $Z_{n-1}$ for each category. When generating high-quality negative example representations for a specific subject category in NICO, we input one sample representation from each of the other categories into the negative example generator. That is, when generating higher-quality negative examples for $z_{ori}$, we require the $\{z_{\text{neg}_{m}}\}_{m=1}^{M}$ selected during the establishment of the representation group in the D3C method to serve as inputs to the negative example generator. Then, the corresponding memory banks $M_{n-1}$ and $Z_{n-1}$ are selected based on the subject category to assist in the generation process. During training, we utilized the Adam[47] optimizer with a learning rate of $10^{-3}$. The Adam optimizer contains a common deep network optimization method. [48]. The experimental results are recorded in Table IV. D4C method generates 5 negative examples for each NICO sample.
@@
TABLE V: Reasoning accuracies on NICO 8:2
@@ IX-B Experiment on RAVEN
We tested Lico-Net on the RAVEN and I-RAVEN datasets to showcase its potential. In our exploration of the maximum potential of Lico-Net in solving RPM problems, we decided to discard the use of Sinkhorn distance for probabilistic representations decoding and instead adopted a Transformer-Encoder[4] equipped with learnable class vectors. Initially, we employed the Sinkhorn distance in preliminary experiments to investigate how Lico-Net induces RPM concepts. However, in order to accelerate model convergence and enhance its flexibility, we have chosen the Transformer-Encoder to undertake the task of representation decoding.
Additionally, we conducted experiments combining D3C and D4C with Lico-Net to showcase the effectiveness of the D3C and D4C methods. The D4C method generates 5 negative examples for each RAVEN instance. For a fair comparison, we used the same settings and equipment as the current state-of-the-art (SOTA) method, Triple-CFN[37], including data volume, optimizer parameters, data augmentation, batch size, and other hyperparameters.
@@
An interesting phenomenon observed in Table VI indicates that existing excellent RPM reasoning engines, like Lico-Net, exhibit varying accuracies on the RAVEN[16] and I-RAVEN[40] datasets. I-RAVEN improves the design of negative examples by increasing the number of attribute modifications and the magnitude of deviations, aiming to rectify the unreasonable designs within the option pool of the RAVEN dataset. However, its reasoning logic and problem-solving path have not undergone fundamental changes, and thus its difficulty has not been substantially increased. Surprisingly, I-RAVEN, which does not change in reasoning difficulty, triggers variations in the performance of reasoning engines—a phenomenon that previous works have noticed but failed to explain. However, this paper elucidates the correlation between the quality of negative examples and the precision of learning concepts by reasoning engines, providing a reasonable explanation for this phenomenon, which constitutes one of the theoretical contributions of this paper.
Table VI further demonstrates that Lico-Net, as a reasoning engine that learns probabilistic representations, achieves significantly superior performance compared to conventional reasoning engines based on deterministic representation learning. Furthermore, D4C not only enhances Lico-Net’s performance but also, to some extent, bridges the gap in Lico-Net’s performance between the RAVEN and I-RAVEN datasets.
@@ IX-C Experiment on Bongard-Logo
We experimented with Lico-Net-Bongard on the Bongard-Logo dataset and explored the enhancement of D3C and D4C algorithms on various Bongard-Logo reasoning engines. D4C generates 5 negative examples. Results are in Table VII. Notably, D4C excels at generating probabilistic representations of high-scoring negative examples for Triple-CFN and Lico-Net-Bongard, but this can cause gradient explosion during training. Our solution is to use a weighted summation of generated and dataset-provided negative example representations, with a 0.2::::0.8 ratio. During training, we utilized the Adam optimizer with a learning rate of $10^{-3}$ and a weight decay of $10^{-4}$.
@@
This paper found that augmentation experiments were conducted on both PMoC [36] and Triple-CFN [37] using techniques like rotation and flipping to expand the Bongard-Logo dataset, without significantly altering concept diversity or proportions [36]. The aim was to explore model performance on a larger dataset. We verified D4C’s effectiveness on the augmented dataset through experiments, with results in Table VIII.
@@
The results in Tables VII and VIII indicate that Lico-Net-Bongard exhibits superior performance on the augmented Bongard-Logo problem. This suggests that Lico-Net-Bongard, designed based on probabilistic representation learning, may possess greater network capacity and stronger learning abilities. Especially when combined with the D4C algorithm, Lico-Net-Bongard’s performance on the augmented Bongard-Logo problem is particularly outstanding, fully demonstrating the effectiveness and potential of probabilistic representation learning as a methodological guide for deep learning algorithms, and verifying the efficacy of the adversarial learning-based negative example quality improvement algorithm designed in this paper.
@@ IX-D Experiment on PGM
In this paper, we conducted experiments related to Lico-Net on PGM, where D4C still generated 5 negative examples, and recorded the results in Table IX. The data recorded in the table reveals many meaningful phenomena.
@@
Firstly, it can be seen from the Table IX that Lico-Net demonstrates superior performance compared to other baseline reasoning engines, which reflects the advantages of learning probabilistic representations of concepts. Secondly, The combinations of Lico-Net+D3C, Triple-CFN+D3C, and RS-Tran+D3C showed limited improvement in reasoning accuracy, possibly due to the fundamental differences among RPM, NICO, and Bongard-Logo. D3C’s approach to improving negative sample quality may be more suitable for addressing concept interference and erroneous coupling issues (preliminary experiments indicate that such issues are particularly prominent in Bongard-Logo and NICO), but RPM problems share a common concept space and attribute recognition mechanism, leading to modest improvements. In contrast, D4C significantly enhanced the performance of these models, supporting our hypothesis that better negative samples can improve the performance of deep learning algorithms. This finding holds significant theoretical and practical importance. Lico-Net has greatly benefited from the support of D4C. We also conducted generalization experiments on Lico-Net, D3C, and D4C, with the results presented in Table X.
@@ IX-E Ablation Study
We removed the linear attention mechanism from the D4C generator while retaining the process of the transformer-encoder for processing input representations, and conducted relevant experiments again. The results are recorded in the entry labeled “D4C-” in the Table XI, XII and XIII.
@@
It can be observed from the experimental results recored in these three tables that the step-wise linear attention mechanism serves as a crucial external memory bank for the D4C negative example generator, providing performance improvements.
@@ X Conclusion
This paper points out that when deep learning is applied to problems that embed complex and deep human concepts, deep learning algorithms should explicitly extract these concepts as probabilistic representations and reason based on these representations. The Lico-Net and Lico-Net-Bongard designed in this paper validate this viewpoint and confirm probabilistic representation learning as a key development direction for deep learning.
More notably, this paper proposes that deep learning models tasked with discriminative tasks are more concerned with the distinctions between positive and negative examples rather than the logic of discriminating between them. Focusing on the distinctions between positive and negative examples is feasible and effective in traditional discriminative tasks. However, when the quality of negative examples is not high enough to strictly define the boundaries of concepts, the performance of the algorithm may be compromised. Therefore, this paper suggests that providing stricter negative examples for deep learning models can effectively improve their performance. Consequently, this paper designs the GSCLM algorithm for NICO data and a negative example generator based on a step-wise linear attention mechanism for RPM problems, Bongard-Logo problems, and NICO problems. The experimental results demonstrate the effectiveness of these algorithms, with the D4C generator representing a novel attempt in probabilistic representation learning, hinting at the potential of probabilistic representation learnings in the field of AI.